OpenStudy (anonymous):

perform operation and write answer in lowest term 4/5q-20 -1/3q-12

OpenStudy (gorv):

\[\frac{ 4 }{ 5*(q-4) } -\frac{ 1 }{ 3*(q-4) }=\frac{ 1 }{ q-4 } *(\frac{ 4 }{ 5 }-\frac{ 1 }{ 3 })\]

OpenStudy (gorv):

\[\frac{ 1 }{ q-4 }*\frac{ 7 }{ 15 }=\frac{ 7 }{ 15*(q-4) }\]

OpenStudy (gorv):

@talita

OpenStudy (anonymous):

thanks